OpenStudy (anonymous):

wat is 9 divided by 0.75

OpenStudy (anonymous):

12

jimthompson5910 (jim_thompson5910):

9/0.75 = 9/(3/4) = 9*(4/3) = 36/3 = 12 So 9/0.75 = 12

OpenStudy (anonymous):

thnx